Circular No. 6564

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 R3 On 1996 Oct. 10 C.-I. Lagerkvist, Uppsala Observatory, reported his

discovery of a comet on exposures obtained with the 1-m Schmidt at the European Southern Observatory a month earlier. Nugatory recovery attempts were made in October and November (limiting mag 20.5–21.0) on the basis of orbit computations by the undersigned. Nevertheless, additional September detections of the comet were subsequently located on images obtained in the course of the Spacewatch and NEAT programs. Available observations:

1996 UT α_{2000} δ_{2000} m_{1} Observer
Sept.12.28682 23^{h}09^{m}50.^{s}82 −0^{o}49′35′′.3 20.7 Scotti
13.15139 23 09 34.78 −0 54 59.2 19.5 Lagerkvist
14.27361 23 09 13.81 −1 02 03.7  _{”}
14.40286 23 09 11.71 −1 03 09.7 19  Helin
14.44208 23 09 10.97 −1 03 14.7  _{”}

J. Scotti (Kitt Peak). 0.9-m Spacewatch telescope. Coma diameter ∼ 17′′,

tail 0′.55 long in p.a. 302^{o}. C.-I. Lagerkvist, S. Mottola and U. Carsenty (European Southern Obs.).

Measurer M. Lundström. Reduced by O. Hernius. 60-min and 30-min exposures. Object diffuse, coma more extended than for P/1996 R2.

E. Helin (Haleakala-NEAT/GEODSS). 1-m Ritchey-Chrétien + CCD. Mea-

surers D. L. Rabinowitz and K. J. Lawrence. Object faint and diffuse.

The following parabolic orbital elements are from four of the above observations. It is quite possible that the comet is of short period.

T = 1996 May 30.292 TT  ω = 99.^{o}606
Ω = 196.029 }2000.0
q = 1.76959 AU  i = 5.011
1996 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Aug. 5 23^{h}17.^{m}73 + 2^{o}22.^{′}2 1.069 1.965 141.^{o}2 18.^{o}9 19.6
15 23 17.81 + 1 53.8 1.077 2.023 150.9 14.1 19.7
25 23 15.78 + 1 05.1 1.104 2.086 161.1  9.0 19.9
Sept. 4 23 12.61 + 0 04.0 1.152 2.154 171.3  4.1 20.1
14 23 09.32 − 1 00.4 1.222 2.226 174.6  2.4 20.4
24 23 06.83 − 1 59.6 1.317 2.301 165.3  6.3 20.7
Oct. 4 23 05.80 − 2 47.2 1.434 2.380 155.2 10.2 21.0
14 23 06.52 − 3 19.8 1.574 2.461 145.4 13.3 21.4
24 23 09.09 − 3 36.3 1.733 2.544 136.0 15.8 21.7
1997 February 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden